Let X be a random variable with the following probability distribution:

x : 3 6 9

P(X=x) : 1/6 1/2 1/3

Find E(X) and E(X^2)and using the laws of expectation, evaluate E . (2X+1)^2.


1
Expert's answer
2021-01-26T04:12:55-0500

"E(X)=\sum{_{i=3}^9}XiP(Xi)"

"=3(\frac{1}{6})+6(\frac{1}{2})+9(\frac{1}{3})"

"=\frac{1}{2}+3+3"

"=6\frac{1}{2}"

"E(X^2)=\sum{_{i=3}^9}Xi^2P(Xi)"

"=9(\frac{1}{6})+36(\frac{1}{2})+81(\frac{1}{3})"

"=\frac{3}{2}+18+27"

"=46\frac{1}{2}"

"E(2X+1)^2=E[(2X+1)(2X+1)]"

"=E(4X^2+4X+1)"

"=4E(X^2)+4E(X)+1"

"=(46\frac{1}{2}×4)+(4×6\frac{1}{2})+1"

"=186+26+1"

"=213"